Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
Paoletti, Fiori, Tringali - 15:27 Thursday 06 April 2023 (59669) Print this report
CEB broadband magnetic noise emission test

Today, taking advantage of a stable lock and with the agreement of the commissioning crew, we injected broadband magnetic noise (10 to 200 Hz) into the CEB from 10:54:00 UTC to 11:43:00 UTC (the interferometer unlocked a few seconds later for unknown reasons).

It is interesting to note high coherence with Hrec_hoft (and somewhat less with LSC_DARM) in the whole range of the noise injection.

here a log x plot, and here a linear x plot  (in both Purple is when injection was OFF).

Images attached to this report
Comments to this report:
paoletti - 10:51 Friday 07 April 2023 (59686) Print this report
Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×